OpenStudy (anonymous):

The volume v of a wooden beam is v=ls^2 where l is the length of the beam and s is the length of one side of its square cross section. If the volume of the beam is 1200 in ^3 and its length is 96 in what is the side length?

OpenStudy (johnweldon1993):

\[\large V = ls^2\] We are given V to be 1200 and l to be 96 So we need just solve for 's' \[\large 1200 = 96\times s^2\] Meaning \[\large s^2 = \frac{1200}{96}\] Or \[\large s = \sqrt{\frac{1200}{96}}\]

OpenStudy (anonymous):

Thank you !!!

OpenStudy (johnweldon1993):

No problem!! :)